OpenStudy (goformit100):

Which of the following is a prime number ? A. 33 B. 81 C. 93 D. 97 mentally ?

OpenStudy (goformit100):

@91004775

OpenStudy (agent0smith):

Again, prime just means only divisible by one and itself. Which number is not divisible by anything other than 1 and itself? try dividing by 2, 3, 4, etc. Note there is a rule for a number being divisible by three - the digits add up to a multiple of three. What do the digits of 33 add up to? 81? 93?

OpenStudy (agent0smith):

For 33: 3+3 = 81: 8+1 = 93: 9+3 =

OpenStudy (unklerhaukus):

all but one of those option is divisible by three

OpenStudy (unklerhaukus):